OpenStudy (anonymous):

A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture? If x represents the pounds of $1.50 chocolates used, then which of the following represents the pounds of 90¢ chocolates used? x 30 30 - x

OpenStudy (anonymous):

Feel free to tag me in questions if you need my help. X = chocolates @ $1.50/pound (30-X) = chocolates @ 0.90/pound X * 1.50 + (30-x) * .90 = 30 (1.10) 1.5X + 27 - .9X = 33 .6X = 6 6X = 60 X = 10 10 pounds of the $1.50 chocolates 20 pounds of the $0.90 chocolates

OpenStudy (anonymous):

......so c

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

kk thanks